### v2

# RAPID-Net: Accurate Pocket Identification for Binding-Site-Agnostic Docking

Abstract
Accurate identification of druggable pockets and their features is essential for structure-based drug design and effective downstream docking. Here, we present RAPID-Net, a deep learning-based algorithm designed for the accurate prediction of binding pockets and seamless integration with docking pipelines. On the PoseBusters benchmark, RAPID-Net–guided AutoDock Vina achieves 54.9% of Top-1 poses with RMSD $<2\,\text{\AA}$ and satisfying the PoseBusters chemical‐validity criterion, compared to 49.1% for DiffBindFR. On the most challenging time split of PoseBusters aiming to assess generalization ability (structures submitted after September 30, 2021), RAPID-Net-guided AutoDock Vina achieves 53.1% of Top-1 poses with RMSD $<2\,\text{\AA}$ and PB-valid, versus 59.5% for AlphaFold 3. Notably, in 92.2% of cases, RAPID-Net-guided Vina samples at least one pose with RMSD $<2\,\text{\AA}$ (regardless of its rank), indicating that pose ranking, rather than sampling, is the primary accuracy bottleneck. The lightweight inference, scalability, and competitive accuracy of RAPID-Net position it as a viable option for large-scale virtual screening campaigns. Across diverse benchmark datasets, RAPID-Net outperforms other pocket prediction tools, including PUResNet and Kalasanty, in both docking accuracy and pocket–ligand intersection rates. Furthermore, we demonstrate the potential of RAPID-Net to accelerate the development of novel therapeutics by highlighting its performance on pharmacologically relevant targets. RAPID-Net accurately identifies distal functional sites, offering new opportunities for allosteric inhibitor design. In the case of the RNA-dependent RNA polymerase of SARS-CoV-2, RAPID-Net uncovers a wider array of potential binding pockets than existing predictors, which typically annotate only the orthosteric pocket and overlook secondary cavities.
RAPID-Net, along with interactive demonstration notebooks, is publicly available at: github.com/BalytskyiJaroslaw/RAPID-Net, offering a user-friendly and fully reproducible framework for pocket-guided docking.
*Corresponding Author: hr6998@wayne.edu
Keywords: Protein–ligand interactions; Blind docking; Binding pocket prediction; Soft mask segmentation; Deep residual network; Convolutional Neural Network;
I Introduction and Problem Statement
Molecular docking is the computational problem of predicting the most favorable ligand binding poses in a protein-ligand complex, given the experimentally determined or computer-simulated protein structure and the initial conformation of the ligand 1. It plays a key role in drug development, typically following target identification when the biological molecule responsible for a disease is identified and validated for therapeutic intervention 2. Once the target is identified, molecular docking is used for structure-based drug discovery and development 3; 4 to identify and optimize potential lead compounds 5; 6. The docking results can guide virtual screening workflows, enabling the selection of promising candidates from vast chemical libraries 7; 8. This process helps to determine whether a compound has potential for drug development 9. By significantly improving the speed and efficiency of early drug discovery, molecular docking reduces the time and costs associated with the development of new therapeutics 10.
Typical protein-ligand docking pipelines rely on users specifying the binding pocket, and docking programs such as AutoDock Vina 11; 12, GOLD 13, and Glide 14 use grids to confine their search to known or hypothesized protein’s interaction sites. However, in the absence of such information – in “blind” or binding-site-agnostic settings – protein-ligand docking becomes significantly more difficult 15, as the docking algorithm must scan the entire protein surface, dramatically increasing the computational complexity. Traditional blind docking methods often rely on extensive sampling to explore potential binding sites across the whole protein, but this approach is computationally and time-intensive, making it impractical for large-scale virtual screening tasks 16; 17. With the advent of advanced protein structure prediction methods 18–including AlphaFold 19, ColabFold 20, OpenFold 21, and RosettaFold 22–the number of protein structures generated has surged, often without any corresponding ligand information. Meanwhile, only about $\sim 5\%$ of the human coding genes currently serve as commercial drug targets, leaving a wide range of disease-related targets unexplored 23. Despite this immense potential, most new drug research and development remain focused on a limited set of well-established targets, underscoring the urgent need for novel druggable targets 24. As more proteins lacking known binding pocket information are considered in drug discovery 25, the demand for ligand docking approaches that do not rely on prior knowledge of the binding site has grown significantly, highlighting the necessity for binding-site-agnostic methods.
In this work, we aim to develop a high‐accuracy pocket-finding algorithm (RAPID-Net) optimized for seamless integration with standard docking pipelines to deliver efficient guidance in binding-site-agnostic workflows. Many protein binding site predictors are benchmarked solely on the geometric properties of their predicted pockets 26, but their impact on the accuracy of downstream ligand docking is rarely assessed. While many factors affect docking performance, we show further in the text that pocket prediction quality has a decisive role in determining docking success. We evaluate RAPID-Net’s performance both by docking accuracy – using the widely adopted AutoDock Vina 12 – and by pocket-ligand intersection rates, as detailed further in the text. Indeed, RAPID-Net can provide the search grid centers and dimensions for any docking engine, and Vina was chosen due to its popularity.
Despite relying on Vina’s relatively simple scoring function, RAPID-Net-guided docking outperforms the state-of-the-art blind-docking tool DiffBindFR 27 by a substantial margin (54.9% vs 49.1%) on the PoseBusters benchmark 28 (see Section VII). This result underscores our central claim: precise pocket identification is a decisive driver of downstream docking success. It is important to note that, unlike DiffBindFR, which incorporates the ligand-induced protein changes, in our approach, we use our pocket predictor to guide AutoDock Vina that docks to a rigid receptor, and all evaluations are performed on holo structures. Nevertheless, the same pocket and search grid localization can boost flexible-receptor engines as well, for example, DSDPFlex 29:“By leveraging prior knowledge or information, this innovative approach is anticipated to enhance the search process within the appropriate binding pocket.”. Overall, our results indicate that accurate pocket localization currently outweighs receptor flexibility as the key bottleneck, making RAPID-Net a practical route to higher accuracy in binding-site-agnostic workflows and a solid foundation on which ligand-induced conformational refinements can later be added as a subleading effect. We refer the treatment of ligand-induced conformational changes to our future work, focusing here on maximizing the pocket prediction accuracy.
For comparison with co-folding approaches such as AlphaFold 3 (AF3) 30, we follow the PoseBench protocol 31. This protocol uses a curated subset of 130 protein-ligand complexes from PoseBusters 28 deposited after September 30, 2021, specifically designed to evaluate generalization performance rather than memorization 31. On this challenging benchmark, RAPID‑Net + Vina achieves a Top-1 success rate of 53.1%, only six percentage points below AF3’s 59.5% (see Section VII), while requiring substantially fewer computational resources and offering significantly better scalability for virtual screening.
RAPID-Net supplies an accurate search box for subsequent lightweight rigid-receptor redocking, whereas AF3 predicts the full protein-ligand complex from the protein sequence and the ligand SMILES. The significantly lower computational demand of our approach is exemplified by the 8F4J structure 28, which AF3 was unable to process as a whole: “Another PDB entry (8F4J) was too large to inference the entire system (over 5,120 tokens), so we included only protein chains within $20\ \rm{\AA}$ of the ligand of interest.”. In contrast, RAPID-Net guided Vina to a pose with $\mathrm{RMSD}<2\,\text{\AA}$ [^1] A demonstration video is available at: https://youtu.be/EkUKmoW11pE?si=3aKBkL3ZRq8ibWqo, as illustrated in Fig. 2. Thus, RAPID-Net provides near-state-of-the-art accuracy while remaining practical for large-scale virtual screening.
In terms of the underlying pocket prediction function, existing pocket predicting algorithms can be broadly classified into classical and Machine Learning (ML)-driven approaches 26. Classical methods rely on expert-defined rules or heuristics to detect pockets, whereas ML-driven algorithms learn to extract features from protein data without explicit human instructions. In this work, we adopt the definition of Machine Learning by François Chollet as “the effort to automate intellectual tasks normally performed by humans” 32. From this perspective, the main objective of an ML framework is to unveil a meaningful data representation that allows pocket-prediction rules to emerge automatically, rather than being manually hardcoded.
Furthermore, “Deep Learning (DL) is a specific subfield of Machine Learning: a new take on learning representations from data that puts an emphasis on learning successive layers of increasingly meaningful representations” 32. In the context of pocket identification in our study, these layers of data representations are realized through Deep Neural Networks (DNNs). Finally, in contrast to DL, so-called “shallow” learning approaches typically employ only one or a small number of consecutive data representation layers 32.
In the following Sections, we describe our DL model for pocket identification, named RAPID-Net (ReLU Activated Pocket Identification for Docking). We tested RAPID-Net on various benchmarks to demonstrate its efficiency and analyzed its predictions on several therapeutically important proteins. In Section II, we provide a brief overview of existing pocket prediction algorithms. In Section III, we discuss the architecture and design rationale behind our model. In Section IV, we present the RAPID-Net training pipeline, highlighting its key differences from existing approaches, including the use of ReLU activation in the last layer of the network to operate on a “soft” rather than binary labels. Using the RAPID-Net model developed in Sections III and IV, in Section V we integrate it into a docking protocol that is subsequently used for docking. In Section VI, we describe the evaluation metrics used to assess the quality of the model’s predictions. Sections VII and VIII report docking results on the PoseBusters 28 and Astex Diverse Set 33 datasets, respectively. For completeness and to provide a direct side-by-side comparison with existing pocket prediction algorithms, Section IX evaluates the geometric characteristics of predicted pockets on the Coach420 34 and BU48 35 datasets. To demonstrate the ability of RAPID-Net to identify distant binding sites, Section X analyzes its predictions for several therapeutically important proteins with known distal sites. Finally, in Section XI, we reach our conclusions and outline the potential directions for future work.
II Overview of the Available Pocket Prediction Algorithms
Pocket prediction methods exploit a variety of different techniques to identify potential binding sites. Geometry-based techniques like Fpocket 36, Ligsite 37 and Surfnet 38 identify cavities by analyzing the geometry of the molecular surface of a protein and usually rely on the use of a grid, gaps, spheres, or tessellation 36; 37; 38; 39; 40; 41; 42; 43. Energy-based methods such as PocketFinder 44 rely on the calculation of interaction energies between the protein and chemical group or probe to identify cavities 44; 45; 46; 47; 48; 49 .
Conservation-based methods make use of sequence evolutionary conservation information to find patterns in multiple sequence alignments and identify conserved key residues for ligand binding site identification 50; 51; 52.
Template-based methods rely on structural information from homologues and the assumption that structurally conserved proteins might bind ligands at a similar location 53; 54; 34; 55; 56; 57. Combined approaches or meta-predictors combine multiple methods, or the use of multiple types of data to infer ligand binding sites, e.g., geometric features with sequence conservation 35; 58; 59; 60; 61; 62; 63.
Although classical geometry-based tools such as Fpocket are fast and widely applicable, they can sometimes merge individual pockets into a single “blurred” cavity, for example, as observed in key SARS-CoV-2 targets 64; 65. In Section X, we revisit the Nsp12 protein considered in 64; 65 using the RAPID-Net model we developed to demonstrate its ability to identify functionally significant sites.
III Design Rationale behind our Model
Figure 1: 8DP2 protein structure from the PoseBusters 28 dataset. The Top-1 Vina pose is docked in a predicted sub-pocket devoid of true ligand binding pose, while the subleading pose correctly occupies a second predicted sub-pocket and passes all validation tests.
Figure 2: For the 8F4J protein structure from the PoseBusters 28 dataset, our docking approach achieves $\text{RMSD}<2$, while AlphaFold 3 30 is unable to process the full structure due to input size limitations. Unlike residue-focused methods such as PUResNet V2 66, which produce complex predictions that are difficult to interpret, our approach offers clear, plug-and-play guidance for docking grid selection, as illustrated in the bottom panel.
Figure 3: Schematic representation of RAPID-Net (ReLU-Activated Pocket Identification for Docking). Key improvements that distinguish RAPID-Net from previous approaches include a ReLU activation in the final layer, the usage of a soft dice loss function, including a single SE-attention block, and removing redundant residual connections.
When predicting likely binding pockets, there is a well-known trade-off between precision and recall. A recent evaluation of pocket prediction methods based on the geometrical properties of their predicted pockets 26 highlights this trade-off. Currently available ML methods such as VN-EGNN 67, GrASP 68, and PUResNet 69 achieve high precision (over 90%), but they systematically predict a small number of pockets, leading to low recall. As 26 points out, generating multiple predictions, some of which may be false positives, is often more useful than potentially missing viable binding sites.
To address this shortcoming and mitigate low recall, we develop RAPID-Net as an ensemble-based model to improve the stability of prediction accuracy and coverage of potential binding sites. RAPID-Net consists of five independently trained model replicas to aggregate results and increase the reliability of its predictions. For subsequent docking, RAPID-Net returns two types of pockets:
  1. 1.
    Majority-voted pockets: Consisting of voxels predicted by at least 3 out of 5 ensemble models, for high-confidence predictions.
  2. 2.
    Minority-reported pockets: Consisting of voxels predicted by at least 1 out of 5 ensemble models, increasing the overall recall.
As demonstrated on diverse benchmark datasets in the following Sections, including PoseBusters 28, Astex Diverse Set 33, Coach420 34, and BU48 35, the proposed ensemble-based approach yields more stable and reliable performance compared to current ML-driven pocket predictors, both considering the majority-voted and minority-reported pockets.
For instance, consider the 8DP2 protein from the PoseBusters 28 dataset, illustrated in Fig. 1. RAPID-Net predicts two majority-voted sub-pockets connected by a link. However, because no ligand binding occurs in one of these sub-pockets, it constitutes a false positive. By contrast, one of the subleading poses binds within the second predicted sub-pocket and passes all tests. This example demonstrates RAPID-Net’s ability to balance accuracy and recall, providing a comprehensive and robust pocket prediction framework for complex docking tasks. Furthermore, as shown in the following Sections, the minority-reported pockets predicted by our model are often shallow pockets corresponding to secondary binding sites, in contrast to the more prominent majority-voted pockets.
Types of outputs from pocket predictors. The type of output produced by pocket predictors can be divided into three main categories.
Kalasanty 70 and PUResNet V1222PUResNet V169 predicts cavities where the ligand is likely to reside, whereas PUResNet V2 66 predicts residues likely to interact with the ligand. For simplicity, we refer to PUResNet V1 as “PUResNet” throughout the text, as in this work we focus on cavities for generating an accurate search grid for subsequent docking. 69 both detect potential ligand-binding pockets using a voxel-based representation, where each voxel is $2\text{\AA}\times 2\text{\AA}\times 2\text{\AA}$ in size.
In contrast, PRANK 71 and DeepPocket 72 reweight the alpha-spheres identified by the rule-based FPocket 73 algorithm. As a result, these methods inherit FPocket’s initial search space and any associated cascading errors. Moreover, alpha-sphere-based approaches may lack sufficient granularity and can struggle to capture subtle variations among binding sites, as previously noted 64; 65.
GrASP 68, IF-SitePred 74, VN-EGNN 67, and PUResNetV2 66 provide predictions of potential ligand interactions at the residue level. Although these predictions can reveal important binding details, they are less suitable for docking workflows that require a well-defined three-dimensional region for accurate ligand placement and orientation. By precisely defining these regions, the computational search space is reduced as docking is restricted to reasonable binding pockets. As illustrated in the bottom panel of Fig. 2, residue-level predictions can be difficult to interpret when defining a reasonable search grid.
Although some studies reweight residues to guide blind docking 75, and others 25 employ voxelized pocket prediction models–such as PUResNet 69–in tandem with AutoDock Vina 12, our work focuses on developing a voxel-based cavity prediction model that seamlessly integrates into the docking pipeline. This voxel-based approach not only provides well-defined search regions but also facilitates a more modular and interpretable workflow compared to less structured docking strategies, as illustrated in Fig. 4 and discussed further in Section V.
Furthermore, as discussed in Section VII, improving the accuracy of pocket identification directly improves docking results. In particular, AutoDock Vina 12, guided by our pocket predictor, outperforms the state-of-the-art DiffBindFR 27 docking tool, which otherwise scans the entire protein surface in “blind” settings.
Our model architecture. Fig. 3 depicts the architecture of our proposed RAPID-Net model, which is similar to U-Net 76 with encoder and decoder branches. However, we implement several notable adaptations.
First, we include residual connections only in the encoder part of the model, as our experiments show that they are highly beneficial there, but lead to overfitting if included in the decoder. This approach differs from Kalasanty 70, which omits residual blocks altogether 77, and PUResNet 69, which uses them throughout the network.
Second, although the standard SE-ResNet 78 architecture typically includes attention blocks at multiple layers, studies on breast-cancer imaging 79 and Raman spectra classification 80 have shown that using a single attention block can be highly effective while mitigating overfitting. Our experiments similarly suggest that in inherently noisy datasets, adding too many attention modules can amplify noise and degrade performance.
As we demonstrate in the following Sections, adopting a moderate level of attention, limiting residual connections, and incorporating a modified loss function significantly improves performance compared to earlier models.
IV Model Training and Inference Pipeline
We trained our model using the sc-PDB 81 dataset, which contains protein structures with annotated binding sites. In this dataset, cavities are defined using VolSite 82, which maps the pharmacophore properties of nearby protein atoms onto a three-dimensional grid. This method assumes that ligand-induced conformational changes remain relatively small 83, treating each cavity atom as a “pseudoatom” to denote an interaction point rather than a physical atom.
Following 69, we used a curated subset of sc-PDB 81 in which redundant protein structures were filtered out based on their Tanimoto index 84. Each protein structure was then placed in a $36\times 36\times 36$ grid, with each voxel corresponding to a $2\text{\AA}\times 2\text{\AA}\times 2\text{\AA}$ unit cell. We extracted 18 features per voxel using the tfbio 85 package to provide a representation of the protein environment for the model training.
In contrast to previous studies, we apply the medical image segmentation practices 86; 87; 88 to perform soft segmentation of ligand-binding pockets, rather than treating it as a strict binary classification problem. In Kalasanty 70 and PUResNet 69, each voxel is considered part of a binding pocket if it contains at least one cavity pseudoatom. The number of pseudoatoms per voxel is then clipped to the range $[0,1]$, and the model is trained with a Dice loss function as part of a binary classification task:
$\Delta_{\text{Dice}}\left(V^{\text{clipped}}_{\text{true}},V^{\text{clipped}}_{\text{predicted}}\right)=1-\frac{2\left\lvert V^{\text{clipped}}_{\text{true}}\cap V^{\text{clipped}}_{\text{predicted}}\right\rvert}{\left\lvert V^{\text{clipped}}_{\text{true}}\right\rvert+\left\lvert V^{\text{clipped}}_{\text{predicted}}\right\rvert},$ (1)
where $V^{\text{clipped}}_{\text{true}}$ and $V^{\text{clipped}}_{\text{predicted}}$ represent the binary masks of the true and predicted pockets, respectively, $\cap$ indicates their intersection, and $\left\lvert V^{\text{clipped}}_{\text{true}}\right\rvert$ and $\left\lvert V^{\text{clipped}}_{\text{predicted}}\right\rvert$ denote the number of occupied voxels in the true and predicted masks, respectively.
Recently, PUResNet V2 66 sought to improve performance by replacing the Dice loss function with a focal loss 89. However, our extensive experiments indicate that adopting a “soft” approach 86; 87; 88 is better suited in this context. In contrast to the binary “yes/no” classification of occupied voxels, the density of pseudoatoms serves as an indicator of the proximity of a voxel to the interior or boundary of a pocket. Higher densities typically occur near polar or charged residues and ligand functional groups, whereas lower densities are often associated with hydrophobic regions.
Drawing inspiration from medical image segmentation methods 86; 87; 88, we replace the sigmoid output in the final layer with a ReLU 90 (hence the name of our model, ReLU Activated Pocket Identification for Docking, RAPID-Net):
$\sigma\left(x\right)=\frac{1}{1+e^{-x}}\;\;\rightarrow\;\;\text{ReLU}\left(x\right)=\max\left(0,x\right)$ (2)
For the model training, we also replace the conventional Dice loss in Eqn. 1 with its “soft” variant:
$\Delta_{\text{Soft Dice},\ \text{L}^{2}}(V_{\text{true}},V_{\text{predicted}})=1-\frac{2\sum_{\text{Grid}}\left(V_{\text{true}}\,V_{\text{predicted}}\right)}{\sum_{\text{Grid}}\left(V_{\text{true}}^{2}+V_{\text{predicted}}^{2}\right)}$ (3)
Several ways of implementing the “soft” Dice loss have been proposed 87; 88, but we found that this simplest version, based on the $\text{L}^{2}$ norm, works best.
To mitigate class imbalance, as the number of non-interacting voxels sharply exceeds the number of interacting ones, we applied class reweighting using scikit-learn 91. During inference, a voxel is classified as part of a binding pocket if the model output exceeds $\texttt{threshold}\,=0.5$. When ensembling five models, we apply morphological closing using the binary_closing function from the scipy.ndimage package 92 to mitigate sparsity in the predicted pocket regions across model replicas.
Unlike previous approaches that use cavity6.mol2 labels 69; 70, we train our model with threshold-less pocket labels (cavityALL.mol2). In cavity6.mol2, annotations are restricted to regions within $6.5\,\text{\AA}$ of the ligand’s heavy atoms 81, potentially overlooking distal functional regions such as allosteric pockets, exosites, or flexible loops. These sites can critically influence drug binding and resistance, making them key therapeutic targets. Secondary binding sites, which often lie beyond the $6.5\,\text{\AA}$ boundary, mediate interactions with larger substrates or cofactors and help position them within the catalytic pocket, as further discussed in Section X. By adopting threshold-less labels, our model identifies both catalytic and more distant sites, thereby improving predictions for proteins that contain such secondary binding regions.
This approach also makes the Distance Center Center (DCC) metric, which was previously used to evaluate Kalasanty 70 and PUResNet 69, less relevant. DCC defines the distance between the geometrical center of the predicted pocket and the true ligand binding pose, but “tunnels” can extend from the primary binding pocket to distal residues that lie far from the ligand but have significant therapeutic implications (see Section X for examples). Furthermore, the threshold-less pocket definition is less sensitive to the ligand size, thereby improving the generalizability of our model.
V Pocket-informed Docking Protocol and its Rationale
For each predicted pocket, we define the center of the search grid as the average of the maximum and minimum $x$, $y$, and $z$ coordinates of all pseudoatoms in the pocket:
$\text{center}\left(x,y,z\right)=\frac{\texttt{max}\left(x,y,z\right)+\texttt{min}\left(x,y,z\right)}{2},$ (4)
where $\texttt{max}\left(x,y,z\right)$ and $\texttt{min}\left(x,y,z\right)$ are the maximum and minimum coordinates, respectively. The grid dimensions are then given by:
$\Delta x,\Delta y,\Delta z=(\texttt{max}\left(x,y,z\right)-\texttt{min}\left(x,y,z\right))+2\cdot\text{threshold}$ (5)
Fig. 4 illustrates this setup for the majority-voted pocket in the ABHD5 protein 93 with a threshold value of $2$.
Figure 4: Majority-voted pocket predicted by RAPID-Net and its corresponding search grid with $2$ threshold for ABHD5 protein 93.
Figure 5: For the 8FAV protein structure, our model predictions are inaccurate, and none of the true ligand binding poses are within one of the predicted pockets. Nevertheless, the docking is successful because the search grids with large thresholds smooths out this drawback.
For each majority-voted pocket, we generate search grids with thresholds of $2\,\text{\AA}$ and $5\,\text{\AA}$. For minority-reported pockets, we use thresholds of $2\,\text{\AA}$, $5\,\text{\AA}$, $10\,\text{\AA}$, and $15\,\text{\AA}$. This strategy accounts for large ligands, such as those in the PoseBusters 28 dataset, which may extend beyond the predicted pockets, and also reduces potential inaccuracies in our model’s pocket predictions.
As an example, Fig. 5 shows the 8FAV protein from the PoseBusters dataset 28, where none of the predicted pockets overlap with one of the true ligand binding poses. However, the docking still succeeds because the expanded search grid thresholds encompass the true binding site. This case demonstrates how geometry-based evaluation metrics, such as those in 26, can overlook the practical success of docking, even if the pocket predictions are imperfect.
To demonstrate that the accuracy of our model is primarily a result of architectural and training improvements–not merely ensembling–we also provide single-model predictions from Kalasanty 70, PUResNet 69, and individual RAPID-Net runs. Similarly, for each predicted pocket, we generate grids with $2\,\text{\AA}$, $5\,\text{\AA}$, $10\,\text{\AA}$, and $15\,\text{\AA}$ thresholds.
Finally, we perform molecular docking on each identified grid using AutoDock Vina 1.2.5 12, with default settings (Exhaustiveness$=32$, num_modes$=40$), following the protocol described in 28.
VI Test Benchmarks and Evaluation Metrics
We evaluate docking performance as the percentage of Top-1 Vina poses with a root mean square deviation (RMSD) below $2$ between the predicted and one of the true ligand poses if multiple true poses are available. RMSD is calculated as:
$\text{RMSD}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\lVert\mathbf{r}_{i}^{\text{true}}-\mathbf{r}_{i}^{\text{pred}}\rVert^{2}},$ (6)
[…]
Docking methods like AutoDock Vina 12 balance speed and accuracy 95, utilizing “fast and dirty” scoring functions. A common strategy to enhance their accuracy is to generate an ensemble of potential poses and then identify the best pose among them using more accurate scoring tools, for example 96; 97. Fast scoring functions enable the efficient exploration of numerous poses, while more accurate methods are reserved for refinement and final pose selection. To account for this approach, we also report the sampling RMSD accuracy, defined as the percentage of cases with “at least one correct pose in the ensemble”. This metric underscores the potential gains that can be achieved through more accurate rescoring. Since our primary focus is on improving pocket identification rather than developing new rescoring functions, we defer such work to the future. Additionally, we evaluate the PB-success rate, defined as the percentage of poses with $\text{RMSD}<2$ that pass all chemical validity checks defined by PoseBusters 28.
Finally, similarly to 69, we compute the Pocket-Ligand Intersection (PLI) score to quantify the proportion of ligand atoms residing within the predicted pocket. Unlike 69, which computes a voxel-based intersection, we measure the average fraction of ligand heavy atoms located within $5$ of at least one pocket pseudoatom, yielding a more intuitive metric:
$\text{PLI}=\frac{N_{\text{within 5 \AA}}}{N_{\text{total}}},$ (7)
where $N_{\text{within 5 \AA}}$ is the number of ligand heavy atoms within $5$ of any pocket pseudoatom, and $N_{\text{total}}$ is the total number of heavy atoms in the ligand.
However, unlike 69, we compute this ratio for all protein-ligand pairs in the dataset rather than restricting it to cases where the ligand is centered within $\leq 4\,\text{\AA}$ of the pocket center. If no pockets are predicted, the PLI is set to zero. When a protein has multiple predicted pockets or multiple ligand poses, we report the maximum PLI value across all possibilities. For completeness, we also indicate how many proteins have at least one predicted pocket and how many have at least one ligand pose within the defined search grids.
Our primary dataset for evaluating docking performance is the PoseBusters 28 dataset, which consists of novel protein-ligand complexes that were not available during the development of current docking programs and thus provides a rigorous assessment of their generalization capabilities. For instance, several docking algorithms that claimed superior accuracy on CASF 2016 95 dataset, perform drastically worse on PoseBusters 28. Additionally, we evaluate docking accuracy on the Astex Diverse Set 33. Although this dataset is older than PoseBusters 28, it remains a challenging benchmark for many algorithms that had previously reported high accuracy on CASF 2016 95.
We distinguish between two docking scenarios: those that use “prior knowledge” of the binding site (which is typically unavailable in real-world applications) and truly “blind” settings that lack this information. In both the PoseBusters 28 and Astex 33 datasets, when the binding site is assumed to be known, the reference ligand coordinates are used to center a $25\,\text{\AA}\times 25\,\text{\AA}\times 25\,\text{\AA}$ bounding box. However, to evaluate RAPID-Net under truly “blind”, binding-site-agnostic conditions, we omit these reference ligands entirely.
In addition to docking accuracy, for completeness, we report the PLI rates for PoseBusters 28 and Astex 33 datasets. For direct comparison with PUResNet 69, we also evaluate PLI for the Coach420 34 and BU48 35 datasets, excluding any structures that appear in the training set as specified in 69. Lastly, to highlight the advantages of our model architecture and training approach, we report docking accuracy for both single RAPID-Net models and their ensembled version. PLI is calculated for individual models, majority-voted pockets, and minority-reported pockets.
VII Evaluation on the PoseBusters Dataset
We obtained the following results by docking the PoseBusters 28 dataset using the aforementioned protocol. Fig. 6 shows the percentage of predictions with $\text{RMSD}<2\,\text{\AA}$ and the corresponding PB-success rates. When guided by the ensembled RAPID-Net, AutoDock Vina 12 achieved 55.8% RMSD-correct predictions and a 54.9% PB-success rate, outperforming DiffBindFR 27, which scored 50.2% and 49.1%, respectively. This result highlights the crucial impact of accurate pocket identification on the subsequent docking accuracy. Our approach uses targeted docking with the widely used AutoDock Vina 12, whereas DiffBindFR 27 scans the entire protein surface.
Figure 6: Comparison of Top-1 Vina accuracies when guided by different pocket identification algorithms, and a DiffBindFR-Smina 27, for PoseBusters 28 dataset.
Figure 7: Distribution of RMSD values for Top-1 Vina predicted poses in the PoseBusters 28 dataset. When multiple true ligand poses are available, RMSD to the closest one is reported.
Figure 8: Performance comparison on the subset of novel PoseBusters structures as suggested in 31. The RAPID-Net-guided AutoDock Vina achieves 53.1% compared to 59.5% of AlphaFold 3, with significantly lower computational costs.
Figure 9: In 7KZ9 protein structure, RAPID-Net predicts majority-voted pockets around both true binding poses, and the Top-1 Vina pose passes all tests.
Figure 10: For the 8BTI protein structure, the Top-1 Vina pose is incorrect, but the subleading pose passes all tests.
Figure 11: For the 7XFA protein structure, neither the Top-1 nor any of the subleading Vina predicted poses pass the RMSD test, despite the predicted pocket covering the ligand with $\text{PLI}=83\%$.
Furthermore, as shown in Fig. 6, docking driven by individual RAPID-Net models–the ensemble components–outperforms PUResNet 69 and Kalasanty 70 in docking accuracy. Ensembling RAPID-Net models further improves docking accuracy by providing more stable results. The distribution of RMSD values for the Top-1 predicted Vina poses is shown in Fig. 7. Note the issue of low recall discussed above. While the RAPID-Net ensemble predicts pockets in all 308 structures, PUResNet 69 and Kalasanty 70 predict pockets in only 278 and 263 protein structures, respectively. Since we evaluate accuracy in “blind” docking settings, cases with no predicted pockets are treated as failures by default.
Not all predicted pockets are positioned correctly. Of the 308 structures where RAPID-Net predicts at least one pocket, 307 contain at least one true ligand binding pose entirely within a search grid, meaning that docking can be successful in 307 out of 308 cases. In the remaining case, docking would inevitably fail due to an inaccurate search grid location. By contrast, PUResNet 69 and Kalasanty 70 have only 265 and 249 structures, respectively, where at least one true ligand pose is located completely within the search grid. This demonstrates that our model not only accurately identifies pockets in favorable protein structures but also provides reliable predictions across the entire dataset.
For comparison, Fig. 8 presents the evaluation on a subset of novel protein-ligand complexes from the PoseBusters benchmark, as defined by31. Compared to the best-performing structure prediction model, AlphaFold 3, as evaluated by PoseBench, RAPID‑Net + Vina achieves a Top-1 success rate of 53.1%, closely approaching AF3’s 59.5%. This result highlights RAPID-Net’s ability to achieve competitive accuracy while offering a substantially lower computational cost.
Figure 12: Solid bars represent Top-1 accuracies. Dashed segments indicate the accuracy achievable when at least one pose in the ensemble is correct–potentially attainable with a more accurate reweighting tool.
Figure 13: For the 7PRI protein structure, the Top-1 Vina pose fails while a subleading one succeeds. In (a), the majority-voted pockets are shown in cyan, and one of the true ligand binding poses largely overlaps with one of them. In (b), minimally-reported pockets are displayed as purple sticks.
Figure 14: In the 7P1F protein structure, two true ligand binding poses exist, both largely within the majority-voted pockets predicted by RAPID-Net. The Top-1 Vina pose is within one of these pockets but is inaccurate, failing the RMSD $<$ 2 Å test. However, a subleading pose occupies a different pocket and passes all tests.
Figure 15: For the 7NUT protein structure, there are two true ligand binding poses, both largely within majority-voted pockets predicted by RAPID-Net. The Top-1 Vina pose failing the RMSD $<$ 2 Å test and a subleading pose passing the test are in the same majority-voted pocket.
Figure 16: In the 7A9E protein structure, RAPID-Net predicts a single majority-voted pocket. The Top-1 Vina pose is located in an incorrect part of that pocket, whereas a subleading pose that passes the RMSD $<$ 2 Å test–aligning with one of the true ligand binding poses–is found in another part of the same pocket.
Figure 17: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the PoseBusters dataset 28.
For the 8BTI protein structure illustrated in Fig. 10, the Top-1 Vina pose fails, but one of the subleading poses passes all tests. In contrast, for the 7XFA protein structure shown in Fig.11, neither the Top-1 nor any of the subleading Vina poses satisfy the RMSD test, with the closest pose in the ensemble having $\text{RMSD}\,=4.36\,\text{\AA}$. This occurs even though the majority-voted pocket predicted by our model covers most of the true ligand binding pose with $\text{PLI}=83\%$.
Similarly to the 8BTI example in Fig. 10, many protein structures have the following pattern: the Top–1 AutoDock Vina pose fails the RMSD $<$ 2 Å criterion, yet a lower-ranked pose meets it. To illustrate this behavior, we report the sampling accuracy, defined as the fraction of cases in which at least one pose in the ensemble achieves RMSD $<2\,\text{\AA}$. Under this metric, as shown in Fig. 12, AutoDock Vina with the prior knowledge of the search box 12 has a success rate of 93.8%.
In comparison, AlphaFold 3, when run with pocket information available, achieves a Top–1 accuracy of 93.2% with RMSD $<2\,\text{\AA}$ 30. We did not compute AlphaFold 3’s sampling accuracy and therefore make no claim of overall superiority. This comparison merely illustrates that the current limitations in docking accuracy stem more from pose ranking than from conformational sampling.
This difference widens in fully “blind” docking settings: AutoDock Vina, guided by a RAPID–Net, achieves the 92.2% of RMSD $<2\,\text{\AA}$ sampling accuracy, whereas AlphaFold 3 achieves 80.5% of RMSD $<2\,\text{\AA}$ Top–1 accuracy. These results suggest that improved re-ranking algorithms could further boost pocket-guided docking workflows. Furthermore, RAPID-Net performs better than PUResNet 69 and Kalasanty 70 in guiding AutoDock Vina 12, as can be observed in Fig. 12.
To better illustrate this pattern, we consider four representative cases in which the Top–1 AutoDock Vina pose fails to reach the RMSD $<2\,\text{\AA}$ threshold, while a pose with a much lower ranking succeeds.
Fig. 13 illustrates the 7PRI protein structure, which contains two true ligand binding poses–one largely covered by the majority-voted pocket predicted by RAPID-Net, and the other by a minority-reported pocket. The Top-1 Vina pose, which fails the RMSD $<$ 2 Å test, is docked in the minority-reported pocket. A subleading pose in the same pocket passes the test, underscoring the importance of minority-reported pockets in occasionally yielding correct predictions.
Fig. 14 shows the 7P1F protein structure, where both true ligand binding poses lie largely in the majority-voted pockets predicted by RAPID-Net. Although the Top-1 Vina is docked in one of these pockets and fails the RMSD $<$ 2 Å test, a subleading pose lands in another majority-voted pocket, closer to the second true ligand pose, and passes the test.
Fig. 15 illustrates the 7NUT protein with two true ligand binding poses, both largely within the majority-voted pockets. Both the Top-1 Vina pose, which failed the RMSD $<$ 2 Å test, and the subleading pose, which passed the test, are located in the same majority-voted pocket near one of the true binding poses.
Finally, in the 7A9E protein structure shown in Fig. 16, RAPID-Net predicts a single majority-voted pocket. The Top-1 Vina pose is docked in the wrong part of that pocket, while a subleading pose successfully passes the RMSD $<$ 2 Å test by aligning with one of the true binding poses in another part of the pocket.
These results illustrate that the AutoDock Vina 12 sampling mechanism is robust and reliably generates near-native binding poses within the ensemble. However, its main limitation is its ranking power: although it reliably generates poses, it struggles to distinguish correct poses from incorrect ones in its ranking.
Finally, Fig. 17 shows the distribution of PLI values for PUResNet 69, Kalasanty 70, and RAPID-Net for this dataset. All five replicas of RAPID-Net outperform both PUResNet 69 and Kalasanty 70. Majority voting stabilizes the PLI predictions, and including minority-reported pockets further improves the PLI.
To highlight the necessity of model ensembling to provide stable pocket predictions, Table I summarizes the level of pocket and ligand overlap and lists:
  • The number of protein structures with at least one predicted pocket (shown in the “Nonzero” column). Since no protein in all our datasets is a true negative, failure to predict any pocket for a given structure is automatically an incorrect prediction.
  • The number of structures with viable search grids encompassing at least one true ligand binding pose (“Within $15\,\text{\AA}$” column), where the docking may succeed.
  • The average Pocket-Ligand Intersection (PLI) value for all protein structures.
Method Nonzero Within 15 Å PLI
Kalasanty 263 249 74.34%
PUResNet 278 265 79.51%
RAPID Run 1 307 298 90.17%
RAPID Run 2 305 289 87.90%
RAPID Run 3 308 296 88.71%
RAPID Run 4 308 300 88.57%
RAPID Run 5 308 292 86.32%
RAPID ensemble 308 307 91.44%/98.09%
Table I: Comparison of pocket predictions across different methods for PoseBusters 28 dataset consisting of 308 protein structures.
For the RAPID-Net model, individual runs show variability in pocket prediction quality. For example, Run 5 predicts pockets for all 308 protein structures but with lower accuracy, resulting in the lowest average PLI of 86.32% among the five runs. In contrast, Run 1 fails to predict pockets for one structure out of 308 but achieves the highest average PLI of 90.17% among five runs. Combining predictions via majority voting among these RAPID-Net runs increases the average PLI to 91.44%. Further incorporating minority-reported pockets boosts the average PLI even higher, to 98.09%. In the final ensemble of pockets predicted by RAPID-Net, all 308 protein structures have at least one predicted pocket and only one structure lacks a search grid covering at least one true ligand binding pose. These observations highlight the importance of combining pocket predictions from multiple runs to achieve reliable results for subsequent docking.
In the next Section VIII, we perform docking on the Astex Diverse Set 33 following a similar procedure as described for the PoseBusters 28 dataset.
VIII Evaluation on the Astex Diverse Set
Figure 18: Comparison of docking accuracies for the Astex Diverse Set 33 when AutoDock Vina 12 is guided by different pocket prediction algorithms.
Figure 19: Distribution of RMSD values for Top-1 Vina predicted poses in the Astex Diverse Set 28. When multiple true ligand poses are available, the RMSD to the closest one is considered.
Figure 20: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the Astex Diverse Set 33.
Figure 21: Docking accuracy of AutoDock Vina 12 for the Astex Diverse Set 33 when guided by different pocket prediction algorithms, comparing Top-1 and ensemble accuracy.
The docking results of AutoDock Vina 12 guided by different pocket-finding algorithms for the Astex Diverse Set 33, consisting of 85 protein structures, are shown in Fig. 18. The corresponding RMSD distribution of the Top-1 predicted Vina poses is presented in Fig. 19, and the associated PLI scores are shown in Fig. 20.
Several striking features emerge from these graphs. First, although each individual Run of our model outperforms the docking accuracy achieved when guiding AutoDock Vina 12 using PUResNet 69 or Kalasanty 70, ensembling these individual Runs does not improve the overall docking accuracy. This happens even though the ensembled version of RAPID-Net achieves a higher ligand-pocket intersection, as shown in Fig. 20 and Table II.
Furthermore, unlike the PoseBusters 28 dataset, when AutoDock Vina 12 is guided by Kalasanty 70 on the Astex Diverse Set 33, it achieves better docking accuracy than when guided by PUResNet 69, even though PUResNet 69 has more structures with at least one true ligand binding pose completely within the search grid (72 vs. 70, respectively).
Method Nonzero Within 15 Å PLI
Kalasanty 73 70 78.72%
PUResNet 76 72 82.36%
RAPID Run 1 85 83 94.83%
RAPID Run 2 85 79 89.00%
RAPID Run 3 85 78 91.09%
RAPID Run 4 85 83 92.61%
RAPID Run 5 85 81 92.47%
RAPID ensemble 85 84 97.15%/98.82%
Table II: Comparison of pocket predictions across different methods for Astex Diverse Set 33 consisting of 85 protein structures.
Figure 22: In the 1G9V protein structure from Astex Diverse Set 33, two true ligand binding poses are observed within the same majority-voted pocket predicted by RAPID-Net. While the Top-1 Vina pose fails the RMSD test, the subleading one passes all validation tests.
Nevertheless, when considering sampling accuracy–achieved either by the Top-1 pose or any subleading pose, as illustrated in Fig. 21–the ensembled version of RAPID-Net outperforms each of its individual Runs: 95.3% versus 92.9%, 87.1%, 89.4%, 90.6%, and 91.8%, respectively. Additionally, PUResNet 69 and Kalasanty 70 both achieve the same sampling accuracy of 78.8%. A comparison of Tables I and II alongside Figs. 12 and 21 shows that RAPID-Net achieves a higher PLI rate on the Astex Diverse Set 33 than on the PoseBusters 28 dataset. However, the Top-1 docking accuracy is higher for PoseBusters 28, while the “at least one correct pose in the ensemble” accuracy is better for the Astex Diverse Set 33. These results further emphasize and illustrate that, in addition to accurate pocket identification, another major bottleneck in the docking process is the accurate reranking of the generated poses.
In the next Section, we compare our model against other available pocket predictors in terms of PLI metrics using their original test datasets to enable a direct side-by-side comparison.
IX Evaluation on Coach420 and BU48 Datasets
For direct comparison with Kalasanty 70 and PUResNet 69, we evaluate the corresponding PLI rates on the Coach420 34 and BU48 35 datasets. Following 69, we exclude protein structures present in the training sets, resulting in 298 and 62 protein-ligand structures for Coach420 34 and BU48 35, respectively.
Method Nonzero Within 15 Å PLI
Kalasanty 273 248 76.41%
PUResNet 280 259 78.37%
RAPID Run 1 296 285 85.68%
RAPID Run 2 298 279 86.48%
RAPID Run 3 293 264 76.78%
RAPID Run 4 298 288 91.39%
RAPID Run 5 298 275 80.59%
RAPID ensemble 298 297 86.75%/95.49%
Table III: Comparison of pocket predictions across different methods for Coach420 34 dataset consisting of 298 protein structures.
Method Nonzero Within 15 Å PLI
Kalasanty 54 53 77.80%
PUResNet 42 39 46.90%
RAPID Run 1 61 60 86.52%
RAPID Run 2 62 62 97.57%
RAPID Run 3 61 60 86.68%
RAPID Run 4 60 58 84.47%
RAPID Run 5 62 43 46.97%
RAPID ensemble 62 62 95.79%/99.82%
Table IV: Comparison of pocket predictions across different methods for BU48 35 dataset consisting of 62 protein structures.
Figure 23: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the Coach420 34 dataset.
Figure 24: Distribution of the maximum PLI scores corresponding to the pockets predicted by RAPID-Net, PUResNet 69, and Kalasanty 70 for the BU48 35 dataset.
Figure 25: For the Nsp12 protein 64, the pockets predicted by Kalasanty 70 and PUResNet 69 – shown as green and orange dots, respectively – and the likely interacting residues predicted by PUResNet V2 66 correspond to the orthosteric remdesivir binding site 64. In contrast, RAPID-Net not only identifies the orthosteric site – corresponding to the majority-voted pocket, shown as cyan sticks – but also successfully detects putative secondary and allosteric pockets reported in previous studies 98; 99, which correspond to the minimally-reported pockets predicted by our model, shown as purple sticks.
Figure 26: Thrombin (RCSB PDB: 1DWC). Unlike the pockets predicted by Kalasanty 70 and PUResNet 69 –shown as green and orange dots, respectively – the majority-voted pocket predicted by our model, shown as cyan sticks, has a distinct bulge extending toward residues associated with Exosite I.
(a) Human Carbonic Anhydrase I (RCSB PDB: 1AZM). The majority-voted pocket predicted by our model reveals a distinct hook-like bulge that wraps around His64 residue, responsible for the protein transfer to the main binding site.
(b) Human Carbonic Anhydrase II (RCSB PDB: 3HS4). PUResNet 69 does not predict any pockets for this protein structure, but Kalasanty 70 predicts one around the main binding site. One majority-voted pocket predicted by our model covers the active site and the other one is located near the histidine cluster responsible for the proton shuttle.
(c) Human Carbonic Anhydrase II (RCSB PDB: 3HS4). Minority-reported pockets predicted by our model not only cover the active site but also have a $\text{PLI}=92\%$ with the two AZM binding poses in shallow pockets on the enzyme’s surface.
(d) Human Carbonic Anhydrase II (RCSB PDB: 3HS4). The interacting residues predicted by PUResNet V2 66 are located only near the active site, with secondary binding sites not being predicted.
Figure 27: Majority-voted pockets predicted by RAPID-Net are represented by cyan sticks, while the minimally-reported pockets are shown by purple sticks. The interacting residues predicted by PUResNet V2 66 are shown in yellow.
Figure 28: Ls-AChBP (RCSB PDB: 2ZJV). The majority-voted pocket predicted by our model has a distinct bridge towards the Thr155-Asp160 region belonging to the F loop. Kalasanty 70 and PUResNet V1 69 do not predict any pockets for this protein structure, while PUResNet V2 66 predicts interacting residues only in the immediate vicinity of the binding pose.
Unlike the original PUResNet 69 paper where PLI values were evaluated only on protein structures where the distance-center-to-center (DCC) between predicted pocket centers and ligand centers was $\leq 4\text{\AA}$, we report the results for all protein structures, over the whole dataset. This comparison seems more appropriate because a rough initial approximation to set up the search grid is often sufficient to achieve successful docking, as illustrated in Fig. 5. Furthermore, as discussed in the next Section, the RAPID-Net model sometimes predicts meaningful “tunnels” or “bridges” to distant binding sites that indirectly influence ligand binding. In such cases, the DCC metric becomes irrelevant since it does not reflect the functional importance of these remote interactions.
Table III and Fig. 23 summarize the results for the Coach420 dataset 34. Except for Run 3, every RAPID-Net run surpasses PUResNet 69 and Kalasanty 70 in terms of an average PLI score. Furthermore, when comparing the number of protein structures in which at least one true ligand binding pose is entirely within the search grid, up to the largest grid with a threshold of $15\ \text{\AA}$, all RAPID-Net Runs outperform both PUResNet 69 and Kalasanty 70. Similarly, Table IV and Fig. 24 present the results for the BU48 dataset 35. Here, all RAPID-Net Runs, except Run 5, outperform both Kalasanty and PUResNet in terms of an average PLI and the number of protein structures where at least one true ligand binding pose is completely contained within the search grids.
Different RAPID-Net Runs have varying performances across the four test datasets, reflecting the inherent element of randomness in the model training. For example, on the PoseBusters 28 dataset, Run 2 has the fewest protein structures with viable predicted search grids, as shown in Table I. In contrast, Run 3 produces the fewest viable grids for the Astex Diverse Set 33 and Coach420 34, as shown in Tables II and III, while Run 5 shows the lowest PLI for BU48 35, as shown in Table IV. Notably, despite having lower coverage on Coach420 34 and BU48 35, Run 5 achieves the highest “at least one correct pose in the ensemble” rate on PoseBusters 28, as shown in Fig. 12.
These results further emphasize the importance of ensembling model predictions to capture all potential binding sites and account for variability in model training. By combining five RAPID-Net models, we mitigate performance defects of individual model Runs, yielding more robust and reliable results. At the same time, as can be observed from all the data that we presented, across all four test datasets, RAPID-Net consistently exhibits stronger generalization ability than both PUResNet 69 and Kalasanty 70.
In addition to docking accuracy and the coverage of true ligand binding poses by predicted pockets, it is crucial to identify relevant distant sites that indirectly influence ligand binding, which is the topic of the next Section.
X Identification of Allosteric Sites, Exosites, and Flexible Regions for Drug Design
As noted earlier, a common problem with Deep Learning-based pocket prediction methods is that they may miss secondary pockets due to the scarcity of labeled training data, while classical methods may produce “blurry” pockets 65. To demonstrate the ability of RAPID-Net to accurately detect both primary and secondary pockets, we focus on the therapeutically relevant case of the Nsp12 protein, an RNA-dependent RNA polymerase (RdRp), which plays a central role in viral RNA replication and is a key target for the development of antiviral drugs against SARS-CoV-2.
RNA-dependent RNA Polymerase (RCSB PDB: 7BV2). As illustrated in Fig. 25, RAPID-Net successfully identifies the orthosteric remdesivir binding site 64 – containing residues 448, 545, 553, 555, 618, 759, 760, and 761 – as its “majority-voted” pocket, consistent with predictions from other pocket prediction methods, PUResNet V1 69, Kalasanty 70, and PUResNet V2 66. In addition to the main binding site, computational studies have proposed allosteric 98 and secondary pockets 99 that can be used for drug development. Unlike other pocket predictors, RAPID-Net also highlights putative allosteric and secondary pockets reported in the literature 98; 99, demonstrating its ability to recover both primary and secondary binding sites.
Additionally, to illustrate RAPID-Net’s ability to identify remote sites of therapeutic interest, we consider four proteins where such sites are well documented.
Thrombin (RCSB PDB: 1DWC). The RCSB PDB entry 1DWC 100 is the crystal structure of human $\alpha$-thrombin in complex with the inhibitor MD-805. Thrombin plays a pivotal role in the coagulation cascade and is therefore a key target in treating acute coronary syndromes 101. As shown in Fig. 26, Kalasanty 70 and PUResNet 69 predict binding pockets largely surrounding the ligand. By contrast, our model, RAPID-Net, detects an additional bulge extending towards residues 71, 73, 75, 76, and 77. These residues belong to the anion-binding Exosite I, which interacts with negatively charged substrates and cofactors such as fibrinogen, thrombomodulin, and COOH-terminal peptide of hirudin 102; 103; 104; 105; 106. Our model’s prediction, a “tunnel” connecting the active site to the Exosite I in Fig. 26, suggests possible long-range interactions, consistent with previous studies of long-range allosteric communication in thrombin 106.
Human Carbonic Anhydrase I (RCSB PDB: 1AZM) and II (RCSB PDB: 3HS4). Human carbonic anhydrase I (hCAI) in complex with a sulfonamide drug (PDB: 1AZM) 107 and hCAII in complex with acetazolamide (AZM) (PDB: 3HS4) 108 are key targets for glaucoma treatment 109 and diuretic therapy 110, with broader potential for treating obesity, cancer, and Alzheimer’s disease 110. In both isoforms, His64 acts as the primary proton shuttle 111; 112; 113.
As shown in Fig. 27(a), while PUResNet V1 69 and Kalasanty 70 predict the primary ligand binding region in 1AZM, our model additionally predicts a hook-shaped bulge wrapping around His64, suggesting an allosteric interaction.
In hCAII (3HS4), a histidine cluster (His3, His4, His10, His15, His17, His64) mediates proton exchange between the active site and the environment 110. Notably, 3HS4 contains two additional AZM molecules in shallow surface pockets–one previously identified 114; 115 and another in a novel binding site 108.
As shown in Fig. 27(b), RAPID-Net identifies two majority-voted pockets: one around the catalytic site and another one near the histidine cluster. As shown in Fig. 27(c), the minority-reported pockets predicted by our model have $\text{PLI}=92\%$ with these secondary binding poses in shallow pockets. This highlights the ability of our model to predict all binding sites in the protein, both primary and secondary.
In comparison, for this protein structure, Kalasanty 70 predicts only the main binding site, PUResNet V1 69 detects none, and PUResNet V2 66 identifies only the interacting residues within the main binding site.
Ls-AChBP (RCSB PDB: 2ZJV). RCSB PDB ID 2ZJV 116 represents Lymnaea stagnalis acetylcholine-binding protein (AChBP) bound to neonicotinoid clothianidin. AChBP serves as a model for nicotinic acetylcholine receptors (nAChRs) and their allosteric transitions 117, providing insights into therapies for neurological disorders such as Alzheimer’s disease, schizophrenia, depression, attention deficit hyperactivity disorder, and tobacco addiction 118; 119; 120.
For the 2ZJV protein structure, neither Kalasanty 70 nor PUResNet V1 69 predict any pockets. However, the majority-voted pockets from our model and the interacting residues of PUResNet V2 66 illustrated in Fig. 28 are predicted around all five true ligand binding positions.
Furthermore, our model’s pockets include distinct bridges toward residues Thr155-Asp160 in the flexible F-loop, suggesting indirect ligand interactions. This finding aligns with the results from 116, who highlighted that induced-fit movements of loop regions, including the F-loop, are essential for ligand recognition by neonicotinoids.
These examples demonstrate our model’s capability to identify regions influencing ligand binding–even without direct ligand contact. By revealing distal regions such as Exosite I in thrombin, His64 in hCA, and remote loop segments in Ls-AChBP, our approach provides key structural insights that can drive allosteric inhibitor design and broader therapeutic innovation.
XI Discussion and Conclusions
As the number of protein structures without known binding sites continues to grow, performing binding-site-agnostic (or “blind”) docking has become crucial for structure-guided drug design. Successful blind docking relies heavily on accurately identifying the search grid where the ligand is likely to bind. However, most existing pocket prediction tools operate irrespectively of docking pipelines and are evaluated using metrics that do not directly correlate with docking success. To address this gap, we developed RAPID-Net, an ML-based pocket identification tool specifically designed for seamless integration with docking pipelines. We tested RAPID-Net’s effectiveness in guiding blind docking using AutoDock Vina 1.2.5 12, but our approach can be easily adapted to any docking software that requires a well-defined search grid.
When guided by RAPID-Net, AutoDock Vina 12 outperforms DiffBindFR 27 by over 5% in blind docking accuracy on the PoseBusters 28 dataset, highlighting the direct relationship between improved pocket identification and increased docking accuracy. Furthermore, RAPID-Net provides precise and compact search grids, enabling the docking of ligands to large proteins such as 8F4J from the PoseBusters 28 dataset, which tools like AlphaFold 3 30 cannot process in whole, highlighting the critical importance of accurate and focused search areas for cost-effective docking.
We found that another major factor limiting docking accuracy is the reranking of generated poses. In addition to precise pocket identification, developing improved reranking tools could further enhance the overall performance of our combined scheme by more effectively selecting the most favorable poses from the generated ensemble.
We attribute the success of RAPID-Net to the following key changes and innovations we introduced:
  • Soft Labeling and ReLU Activation: Unlike conventional binary segmentation tasks, we applied soft labels and ReLU activation in the output layer, drawing inspiration from medical image segmentation techniques. This approach enables the model to more effectively differentiate between the internal regions of pockets and their boundaries.
  • Attention Mechanism: We integrated a single attention block within the encoder-decoder bottleneck, enhancing the model’s ability to focus on relevant features while reducing the risk of overfit in inherently noisy data environments.
  • Simplified Architecture: By eliminating excessive residual connections, we streamlined the model architecture, resulting in improved performance. This simplification was effective given the noisy nature of the dataset.
Furthermore, RAPID-Net demonstrates the ability to identify “bridges” to distal sites further than $6.5\ \text{\AA}$ from the primary pocket that are often critical for therapeutic intervention. This makes RAPID-Net a promising tool for targeting remote therapeutic sites and broadening the scope of structure-based drug design.
Additionally, beyond improved docking accuracy, in our future work, we plan to integrate RAPID-Net into de novo pocket-conditioned drug design pipelines by guiding the generation of novel compounds based on the shape features of the predicted pockets. Although our current study does not directly exploit these features, using the predicted region solely as a search box, as illustrated in Fig. 4, they may prove highly valuable for generative models 121; 122; 123; 124; 125.
Usage of Our Model and Reproducibility of Results
All code and relevant data (proteins, pockets, poses) are publicly available on GitHub  [^3] https://github.com/BalytskyiJaroslaw/RAPID-Net. To reproduce our results, interested readers should download these files and follow the instructions provided in our repository.
To apply our trained model to predict pockets for other proteins and perform subsequent docking with AutoDock Vina 12, we provide an accompanying notebook with step-by-step guidance in the same repository. Additionally, our pocket predictor can be integrated with other docking tools, for example, by using spherical search grids. Users may also refine the ensemble of poses generated by AutoDock Vina 12 guided by our pocket predictor by any preferred reweighting method, leveraging our pocket predictions to improve docking accuracy.
Author Contributions
YB designed the neural network, conducted model training and docking experiments, and drafted the manuscript. IH and AB contributed to data analysis and provided medical expertise. CVK supervised the research and contributed to data analysis. All authors contributed to the proofreading and approved the final version of the manuscript.
Acknowledgments
Research reported in this publication was supported by the National Institute of Diabetes and Digestive and Kidney Diseases of the National Institutes of Health under award number R01DK076629. We are grateful to Richard J. Barber and the Barber Integrative Metabolic Research Program for financial support. We also thank James Granneman and Kelly McNear for valuable discussions.